# The role of quantum and classical correlations in shrinking algorithms for optimization

### v1

@@ III-E Better quantum correlations lead to improved performance
Fig. 5 showcases the results of our experiments, where the approximation ratio is computed with respect to the Gurobi solutions as defined in Eq. (19). We note that in this case Gurobi was able to certify the optimality of the solutions. For comparison, we also plot the results obtained via LP and SDP correlations. In the figure the heights of the bars correspond to the median approximation ratio across all problem instances and runs of an algorithm. The upper and lower error bars represent the third and first quartiles of the approximation ratios obtained, respectively. We run the shrinking algorithm multiple times for each instance due to the random tie-breaking when choosing a correlation in a shrinking step. This can lead to different final solutions. This effect is particularly important to QAOA, where at low depths $p$ the correlations depend only on the local neighborhood of a given edge. For 3-regular graphs, there is a limited number of distinct local neighborhoods, which leads to many ties in the correlations [30, 49].
Crucially, results in Fig. 5 confirm the intuition that deeper QAOA circuits should generate better correlations. This is evident from the increased performance of the bare algorithms as the depth $p$ increases. Moreover, this steady performance increase transfers over to the performance of the shrinking algorithm, irrespective of the amount of recalculations performed. Furthermore, we again observe that the performance of the shrinking algorithm improves when recalculating the correlations after every shrinking step. In fact, we are able to optimally solve all of the considered graphs already at $p=2$ if the recalculation of correlations is performed after each shrinking step. Nevertheless, it is reasonable to expect that for larger instances higher depths would be required. Finally, we note that the shrinking algorithm with no recalculations ($r=\infty$) on average outperforms the bare algorithm, even at higher depths.
Because of the low density of the 3-regular graphs, the problem instances are in the density regime where the shrinking algorithm of LP and its classical counterpart have the best performance of the different algorithms, as we have seen in section III-C. Thus, all variants of LP reach the optimal solutions. Furthermore, the size of the problems is small enough that also all three SDP algorithms attain very good approximation ratios such that the median approximation ratios are at $100\%$. However, the SDP shrinking variant with $r=\infty$ solves more problem instances perfectly than the bare algorithm. Only the shrinking version with $r=1$ is able to achieve optimal results for all MaxCut instances.
@@ IV Discussion
In this work, we have extended and analyzed algorithms that solve a combinatorial optimization problem by recursively shrinking it. We compared the utility of different methods for computing correlations, both quantum and classical, and thus provided classical benchmarks that quantum-informed shrinking will need to surpass in order to be useful in practice. We remark that our proposed shrinking procedure can be regarded as a standalone heuristic that improves the performance of the bare underlying algorithm. We ran numerical simulations for the shrinking procedure applied to the MaxCut problem.
@@
Interestingly, we observe that the shrinking algorithm not only increases the performance of the quantum algorithm QAOA, as already shown in [7, 11], but can also significantly improve upon the performance of classical algorithms like the Goemans-Williamson algorithm or the linear programming relaxation of MaxCut. Furthermore, the proposed shrinking procedure can serve as a rounding heuristic for the relaxed LP or SDP. Similarly to the spanning tree heuristic, it rounds perfect correlations to the optimal solution. However, our studies indicate that it delivers better integer solutions for fractional correlations than spanning tree rounding.
@@
Applying the shrinking algorithm to Erdős-Rényi graphs of various densities shows that the shrinking algorithm employing LP correlations perform extraordinarily well for low densities, whereas the performance quickly plummets as the density increases. Hence, this suggests a very simple decision metric about the utility of the LP correlations: For solving low-density instances, the LP correlations should be used, but for higher-density instances, other correlations are expected to achieve better results. This sensitivity of the LP to the problem density can be explained by the definition of the relaxation model. Whereas the other relaxations use node-assignment as the decision variables, LP uses decision variables based on edges. This means that the number of variables for a constant number of nodes in the LP increases significantly when increasing the density of the problem graph.
@@
The SDP and the QAOA correlations have a more constant performance where the approximation ratio slightly increases with increasing density. As mentioned before, this can be caused both by an increase in the quality of the shrinking algorithm solution or a decrease in the performance of the reference Gurobi solution.
@@
Furthermore, for all three methods of computing correlations, we observe that more recalculations tend to yield better results. However, this improved performance comes at the cost of an increase in the computational resources required to run the algorithm. Notably, the performance of the LP and SDP-informed shrinking algorithm only improves slightly when recalculating correlations in every step $(r=1)$ compared to every ten steps $(r=10)$.
@@
The correlations from the GW-rounding perform best over the entire range of problem instances considered in this work. Interestingly, while there is still a trend that the recalculations improve the performance, the best results with GW correlations are returned for a recalculation interval of $r=10$. In contrast, for the SDP correlations the performance increases monotonously with the number of recalculations, even though in both cases the correlations are based on the same algorithm. Intuitively, the peak in performance at an intermediate value of $r$ can be understood in terms of the competing effects of recalculations of correlations and GW rounding. Our results generally indicate that increasing the number of recalculations improves the performance. However, performing a recalculation of correlations also requires a new choice of the rounding hyperplane. This new choice of the hyperplane might cause inconsistencies with previously performed shrinking steps. We believe that this trade-off between the amount of recalculations and new choices of hyperplanes is responsible for the observed best performance at $r=10$.
@@
Furthermore, we have confirmed that better quantum correlations improve the performance of the shrinking algorithm by simulating higher-depth QAOA for the correlations used in the algorithm. This result indicates that further improvement of quantum hardware will enable performance gains for shrinking algorithms, thereby providing clues for whether quantum algorithms can provide practically relevant methods for combinatorial optimization.
@@ V Outlook
This study explores the role and practicality of classical and quantum correlations for a recursive shrinking algorithm applied to MaxCut  instances. A natural extension of our work would be to consider other optimization problems. The focus of our paper is primarily the comparison of different sources correlations and their potential use. However, there are several avenues towards increasing the algorithm’s performance.
As discussed in [14], one way to improve the performance of the shrinking algorithm is to not shrink completely until it becomes trivially solvable. Instead, the algorithm should shrink the problem to a size where it can be solved to optimality by a different exact solver. By doing so, we can avoid mistakes during the last few shrinking steps, leading to an improved solution. Moreover, one could envision devising better update rules or using backtracking techniques to rectify shrinking steps that caused a decrease in performance [12].
Furthermore, an interesting property of the graph shrinking algorithm is that with each shrinking step both the graph connectivity and its edge weights change. This is particularly interesting when considering persistency checks for MaxCut, as it is sometimes possible to make assumptions about whether an edge is cut or not by simply considering the MaxCut  graph itself [26, 50]. As the graph changes with each shrinking step, it is possible to conduct these kinds of checks at every shrinking step, even before calculating the correlations, and making a decision for the next step with certainty based on these checks. This can also only increase the performance of the algorithm, but, of course, the checks need to be executed efficiently. On a high-level, this can be seen as an example of a problem-specific shrinking step, that leverages the specifics of the optimization problem [12].
Another way of increasing the performance is the possibility of updating the correlations when shrinking steps without recalculation are performed. For example, if two edges are combined during a shrinking step, the new correlation assigned to this edge could be taken to be e.g., the mean of the correlations belonging to the original edges.
Furthermore, there are numerous other means of obtaining correlations beyond the ones discussed here. For instance, much interest in quantum approaches to combinatorial optimization has been devoted to analog devices [51, 52]. Previously, shrinking algorithms using correlations from an analog device have been proposed in Ref. [12]. We believe that a systematic study of the performance of shrinking algorithms using quantum correlations from analog and gate-based quantum hardware compared to classical benchmarks could provide a useful stepping stone towards a better understanding of the utility of quantum computers for solving COPs.
@@ Acknowledgments
FW thanks Frauke Liers for many fruitful and encouraging discussions. JRF acknowledges Libor Caha and Aron Kerschbaumer for insightful discussions. VF and MP thank Johannes Frank for inspiring dialogues. JRF and MP thank the BMW team for their support. VF, FW, LP and CM are supported by the German Federal Ministry for Economic Affairs and Climate Action (BMWK), project QuaST.

### v2

@@ III-E Better quantum correlations lead to improved performance
Fig. 5 visualizes the results of our experiments, where the approximation ratio is computed with respect to the Gurobi solutions as defined in Eq. (19). We note that in this experiment Gurobi was able to certify the optimality of the solutions. For comparison, we also plot the results obtained via LP and SDP correlations. We run the shrinking algorithm multiple times for each instance due to the random tie-breaking when choosing a correlation in a shrinking step. This can lead to different final solutions. This effect is particularly important for low depth QAOA, where the correlations depend only on the local neighborhood of a given edge. For 3-regular graphs, there is a limited number of distinct local neighborhoods, which leads to many ties in the correlations [30, 51].
Crucially, the results of Fig. 5 confirm our intuition that deeper QAOA circuits generate better correlations. This intuition originates from the fact that the bare QAOA algorithm returns improved solutions as the depth $p$ increases. Fig. 5 shows that the increased performance of bare QAOA for higher depths carries over to the shrinking algorithm, irrespective of the amount of recalculations performed. Furthermore, we again observe that the performance of the shrinking algorithm improves significantly when recalculating the correlations after every shrinking step. In fact, we are able to solve all of the considered graphs to optimality already at $p=2$ if the recalculation of correlations is performed after each shrinking step (yellow bars in Fig. 5). Nevertheless, we expect that for larger instances higher depths are required to obtain optimal solutions. Finally, we note that the shrinking algorithm with no recalculations ($r=\infty$) on average outperforms the bare algorithm, even at higher depths.
The 3-regular graphs have a density of $d=0.06$. Thus, they correspond to the density regime where the LP based shrinking algorithm shows the best performance among the different algorithms (see Fig. 3). Thus, all variants of LP return optimal solutions. Furthermore, due to the smaller size of the problems the SDP algorithms likewise mostly return optimal solutions. The SDP shrinking variant with $r=\infty$ solves more problem instances to optimality than the bare algorithm. However, only the shrinking version with $r=1$ achieves optimal results for all MaxCut instances.
@@ IV Discussion
In this work, we have extended and analyzed algorithms that solve a combinatorial optimization problem by recursively shrinking it. We compared different methods for computing correlations, both quantum and classical, and thus provided classical benchmarks that quantum-informed shrinking algorithms will need to surpass in order to be useful in practice. Importantly, our proposed shrinking procedure can be regarded as a standalone heuristic that improves the performance of the bare underlying algorithm. We evaluated the algorithms through an extensive numerical analysis.
@@
Interestingly, we observe that the shrinking algorithm not only increases the performance of QAOA, as already shown in [7, 11], but can also significantly improve upon the performance of classical algorithms like the Goemans-Williamson algorithm or the linear programming relaxation of MaxCut. Furthermore, the proposed shrinking procedure can serve as a heuristic for rounding LP or SDP relaxations to MaxCut solutions. Our study indicates that the proposed algorithm delivers improved solutions compared to the well-known spanning-tree heuristic.
@@
Applying the shrinking algorithm to Erdős-Rényi graphs of different densities shows that the LP correlations perform very well for low densities, with a rapid decrease in performance as the density increases. This sensitivity of LP to the problem density can be explained by the relaxation model. LP uses decision variables based on edges whereas the other relaxations use node-assignments as decision variables. As a result, the number of variables for a constant number of nodes in the LP increases proportionally to the density of the problem graph. For the other relaxations, the number of variables is independent of the density.
@@
The SDP and the QAOA correlations have a more constant performance where the approximation ratio slightly increases with increasing density. As discussed in Section III, this can be caused by both an increase in the quality of the shrinking algorithm solution or a decrease in the performance of the reference solution.
@@
Furthermore, we observed that more recalculations tend to yield better results. However, this improved performance comes at the cost of an increase in the computational resources required to run the algorithm. Nevertheless, the overhead is at most linear even for recalculation interval $r=1$ Notably, the performance of the LP and SDP-informed shrinking algorithm only improves slightly when recalculating correlations in every step $(r=1)$ compared to every ten steps $(r=10)$.
@@
The correlations from the GW-rounding perform best over the entire range of problem instances. Interestingly, while there is a trend that the recalculations improve the performance, the best results with GW correlations are returned for a recalculation interval of $r=10$. In contrast, for the SDP correlations the performance increases monotonously with the number of recalculations, even though in both cases the correlations are based on the same algorithm. Intuitively, the peak in performance at an intermediate value of $r=10$ can be understood by the competing effects of recalculations and GW rounding. Our results indicate that increasing the number of recalculations improves the performance in general. However, recalculating correlations also requires a new rounding hyperplane. This new hyperplane might cause inconsistencies with previously performed shrinking steps. We believe that this trade-off between the amount of recalculations and new hyperplanes is responsible for the observed best performance at $r=10$.
@@
Furthermore, we simulated higher-depth QAOA with up to three layers for three-regular graphs. Our results confirmed that better quantum correlations improve the performance of the shrinking algorithm. This indicates quantum hardware needs to improve before quantum algorithms can provide valuable approaches to combinatorial optimization.
@@ V Outlook
This study compares classical and quantum correlations for a recursive shrinking algorithm applied to MaxCut  instances. Furthermore, we introduced a new GW-inspired approximation algorithm. To obtain quantum correlations, we used QAOA as a proof-of-concept quantum subroutine. However, QAOA can be readily replaced by other quantum routines for computing variable correlations. A further extension of our work would be to consider other optimization problems. Moreover, there are several avenues towards increasing the algorithm’s performance.
As discussed in [14], one way to improve the performance of the shrinking algorithm is to avoid shrinking the problem until it becomes trivially to solve. Instead, the algorithm shrinks the problem to a size where it can be solved to optimality by an exact solver. This avoids mistakes during the last shrinking steps, leading to an improved solution. Moreover, one could derive improved update rules or use backtracking techniques to rectify shrinking steps that caused a decrease in performance as proposed in [12].
Furthermore, an interesting property of the graph shrinking algorithm is that with each shrinking step both the graph connectivity and its edge weights change. This is particularly interesting when considering persistency checks for MaxCut , which find provably optimal variable assignments [24, 52]. As the graph changes with each shrinking step, it is possible to conduct these checks at every shrinking step, even before calculating the correlations. This can also only increase the performance of the algorithm if the checks are computed efficiently.
Another way of increasing the performance is updating the correlations if shrinking steps without recalculation are performed. For example, if two edges are combined during a shrinking step, the new correlation assigned to this edge could be the mean of the correlations belonging to the original edges.
Furthermore, there are numerous other means of obtaining correlations beyond the ones discussed here. For instance, much interest in quantum approaches to combinatorial optimization has been devoted to analog devices [53, 54], oftentimes paired with a classical optimizer to design the analog protocols [55, 56]. Previously, shrinking algorithms using correlations from an analog device have been proposed in Ref. [12]. We believe that a systematic study of the performance of shrinking algorithms using quantum correlations from analog and gate-based quantum hardware compared to classical benchmarks could provide a useful stepping stone towards a better understanding of the utility of quantum computers for solving COPs.
@@ Acknowledgments
FW thanks Frauke Liers for fruitful discussions. JRF acknowledges Libor Caha and Aron Kerschbaumer for insightful discussions. VF and MP thank Johannes Frank for inspiring dialogues. JRF and MP thank the BMW team for their support. VF, FW, LP and CM are supported by the German Federal Ministry for Economic Affairs and Climate Action (BMWK), project QuaST.